How many grams of O2 gas would you need to make 100grams of water?(H2O)

How many grams of H2 would you need to make 100grams of water?
1
Expert's answer
2020-04-10T13:35:46-0400

The equation for the reaction of the formation of water from oxygen and hydrogen:

2H2 + O2 "\to" 2H2O

  • According to the equation, we can determine the amounts of substances of H2 and O2:

"\nu"(H2O) = "m\/M" = 100(g) / 18(g/mol) = 5.56 mol = "\nu"(H2)

"\nu" (O2) = "\nu" (H2O) "\u2022 1\/2" = 5.56(mol) • 1/2 = 2.78(mol)

  • Then, we can calculate the masses of hydrogen and oxygen:

m(O2) = "\nu • M" = 2.78(mol) • 32(g/mol) = 88.96 g

m(H2) = "\nu•M =" 5.56(mol) • 2(g/mol) = 11.12 g

Answer: m(O2)=88.96 g; m(H2) = 11.12 g